OpenStudy (anonymous):

lim (1/a+h - 1/a) / h , a not equal to 0 h->0 pls answer ! :)

OpenStudy (anonymous):

is it \[\lim_{h \rightarrow 0 }\frac{ 1 }{ a+h }-\frac{ 1 }{ a }\]

OpenStudy (raden):

it means same to find ferivative of f(a) = 1/a = a^-1 so, f'(a) = -a^(-1-1) = -a^-2 = -1/a^2

OpenStudy (raden):

*derivative

OpenStudy (raden):

but if u want using by limit method : lim (h->0) (1/a+h - 1/a)/h = lim (h->0) ((a-(a+h))/(a(a+h)h) = lim (h->0) -h/(a(a+h)h) = lim (h->0) -1/(a(a+h)) just plug h=0; = -1/(a(a+0)) = -1/a*a = -1/a^2

OpenStudy (raden):

if u not sure, ask the god of math "wolfram" http://www.wolframalpha.com/input/?i=lim+%28%281%2F%28a%2Bh%29%29+-+%281%2Fa%29%29%2Fh%2C+h%3D0%29